# Inflationary Gravitational Wave Spectral Shapes as test for Low-Scale Leptogenesis

### v1

@@ II.2 CP Asymmetry
The CP asymmetry $\epsilon_{1}$ is determined by the interference between the tree level and 1-loop diagrams  Flanz:1994yx ; Covi:1996wh and can be calculated from $lH\rightarrow lH$ scattering processes Buchmuller:1997yu . The relevant Feynmann diagrams giving tree level, vertex and self-energy contributions are shown in Fig. 1.
Figure 1: Feynmann diagrams of $N\rightarrow lH$ decay process at the (a) tree-level and (b)-(c) 1-loop level. (b) represents the vertext contribution and (c) represents self-energy contribution.
The maximal asymmetry is obtained for $M_{2,3}\gg M_{1}$ and light neutrino masses $m_{1}\sim\tilde{m}_{1}$, $m_{3}=\sqrt{m_{\text{atm}}^{2}+m_{1}^{2}}$  Buchmuller:2003gz ; Hahn-Woernle:2008tsk ; Buchmuller:2004tu ; Davidson:2002qv ; Hambye:2003rt ,
@@ II.4 Resonant Leptogenesis
Thermal leptogenesis operates at high temperatures, which in turn pushes the inflationary reheating temperature $T_{\rm RH}$ to a high scale, as shown in Eq. (16) and (17). Such high scales involving RHN masses $M_{i}\gtrsim 10^{9}$ GeV are difficult to test in the lab. Also a high neutrino mass is in conflict with Electroweak naturalness condition which requires RHN masses $M_{i}\lesssim 10^{7}$ GeV Vissani:1997ys ; Clarke:2015gwa ; Datta:2022tab . A lower reheating temperature from inflation $T_{\rm RH}$ is motivated in different inflationary scenarios like Drees:2021wgd ; Ghoshal:2022zwu . Even in some SUGRA models Kawasaki:2004qu , reheating temperature is found to be as low as $10^{6}$ GeV which contradicts a high $T_{\text{RH}}$, favouring low-scale-leptogenesis. Motivated by these reasons, and for lab testability, the scale of leptogenesis can be brought down by several orders in the so called resonant regime which focuses on nearly degenerate heavy neutrinos leading to pronounced CP violation in their decays. It can be shown that if the mass difference between the two lightest right-handed neutrinos is of the order of the decay widths of them, i.e. $M_{2}-M_{1}\sim\Gamma_{N_{2}}^{rf}$ the CP parameter $\epsilon_{1}$ is enhanced via the self-energy contribution Pilaftsis:1997dr ; Hambye:2003rt in the loop level diagrams shown in Fig 1.
@@
where, $\epsilon_{1}^{\rm max}$ is the maximal asymmetry in non-resonant case given by Eq. (15). It was shown that within the resonant regime the lowest RHN mass $M_{1}$ can be as low as 1 TeV Pilaftsis:1997jf . In Pilaftsis:2005rv this scale is shown to reach upto the electroweak symmetry breaking scale $\sim 100$ GeV, below which sphaleron processes are suppressed and the lepton asymmetry can not be effectively converted to baryon asymmetry. From here onwards we will consider resonant leptogenesis with $M_{1}\gtrsim 100$ GeV for our analysis. Note that in some alternate leptogenesis scenarios, e.g. via neutrino oscillations, BAU can be generated for even for RHN mass values $\sim 100$ MeV Klaric:2020phc .
III Non-thermal Leptogenesis via scalar decay
In vanilla leptogenesis, RHNs are in thermal equilibrium at production. However, it is possible to produce RHNs non-thermally from inflaton Asaka:1999yd ; Giudice:2003jh ; Hahn-Woernle:2008tsk ; Zhang:2023oyo or a generic scalar decay Datta:2022tab ; Nemevsek:2023yjl . Such a scenario can lead to excess initial abundance of RHNs compared to their thermal abundance. As a result the final efficiency factor, $\kappa_{f}$ can be enhanced upto $\mathcal{O}(2)$ which relaxes the Davidson-Ibarra bound on RHN mass given in Eq. 16. Note that the RHNs can also be produced thermally from the scalar which is similar to thermal leptogenesis. We consider the Lagrangian Ghoshal:2022fud ,
@@ III Non-thermal Leptogenesis via scalar decay
where $f$ represents a Dirac fermion which is in thermal equilibrium with the radiation bath. This kind of Lagrangian can be obtained by considering for example a gauged $U(1)$ extension Ghoshal:2022fud ; Ghoshal:2022hyc of the SM where the vector-like fermions $f$ can couple to the SM sector via the $U(1)$ gauge boson. We will assume the scalar mass to always exceed twice the right-handed neutrino masses, $M_{\phi}>2M_{1}$ where $M_{1}$ is the mass of the lightest RH neutrino. The decay width of $\phi$ to RH neutrinos is Croon:2019dfw ; Ghoshal:2022fud ,
$\Gamma_{\phi\to N_{i}N_{i}}\simeq\frac{y_{N_{i}}^{2}M_{\phi}}{16\pi}\sqrt{1-\frac{4M_{i}^{2}}{M_{\phi}^{2}}}$ (20)
where $i=1,2,3$ represents the 3 RH neutrinos. On the other hand, decay width of the scalar to radiation is given by Ghoshal:2023phi ; Ghoshal:2022fud ,
$\Gamma_{\phi\to R}$ $\simeq$ $\frac{y_{R}^{2}\,M_{\phi}}{8\pi}$ (21)
where subscript $\phi\rightarrow R$ denotes $\phi$ decaying to radiation via the Dirac fermions $f$ in Eq. (19) which are in thermal equilibrium with the SM bath. We define the branching ratios,
$\text{BR}(\phi\to{N_{i}N_{i}})\equiv\frac{\Gamma_{\phi\to{N_{i}N_{i}}}}{\Gamma_{\phi}},\quad\text{BR}(\phi\to{R})\equiv\frac{\Gamma_{\phi\to{R}}}{\Gamma_{\phi}}$ (22)
where $\Gamma_{\phi}=\sum_{i}\Gamma_{\phi\rightarrow N_{i}N_{i}}+\Gamma_{\phi\rightarrow R}$ is the total decay width of $\phi$. Essentially, ${\rm BR}(\phi\to{N_{i}N_{i}})$ is dependent on the coupling $y_{R}$, so decreasing $y_{R}$ will increase ${\rm BR}(\phi\to{N_{i}N_{i}})$. If $y_{R}\sim 0$ then $BR(\phi\to{N_{i}N_{i}})\sim 1$ which leads to maximum non-thermal RHN production from $\phi$ hence maximum lepton asymmetry.
We assume a resonant leptogenesis scenario with $M_{1}\sim M_{2}\ll M_{3}$ such that the scalar dominantly decays to the two lightest RHNs, $N_{1}$ and $N_{2}$. The resonance effect makes sure that enough asymmetry is generated for all $M_{1}$ values starting from 100 GeV, to explain the BAU. RHNs can be in thermal equilibrium with the radiation bath either via their Yukawa couplings $\lambda$ or via the 2-to-2 scattering process $N_{i}N_{i}\leftrightarrow\overline{f}f$ via $\phi$ for non-vanishing values of $y_{R}$. See appendix B for more details. During this paper, we take $y_{R}=0$ when discussing non-thermal leptogenesis and only consider non-zero $y_{R}$ when demonstrating thermal leptogenesis. We treat the independent variables for our analysis to be $M_{i}$, $M_{\phi}$, $\lambda$ and $y_{N_{i}}$. Later in Sec. III we will parameterize $\lambda$ and $y_{N_{i}}$ in terms of the decay temperatures of the scalar $T_{\phi}$ and right-handed neutrinos $T_{N_{i}}$.
@@
We consider the parameter space of the scalar where it dominates the energy budget of the Universe before decaying, leading to a phase of matter dominated pre-BBN era. The RHNs produced from scalar decay, as we shall see, may or may not further dominate the energy density depending on the parameters used. Their out-of-equilibrium decay to relativistic SM particles will inject considerable amount of entropy into the Universe. This dampens the primordial gravitational waves from inflation that re-enters the horizon during the early matter domination epoch. We show such signature of low-scale leptogenesis in primordial gravitational waves can be detected in various GW experiments such as BBO, DECIGO, $\mu$-ARES, ET, LISA, CE etc.
@@ III.1 Set Up
We consider the scalar $\phi$ was in thermal equilibrium with the primordial plasma after the end of inflation either via its direct coupling with SM particles, or via indirect interactions, e.g. coupling with inflaton field in a specific inflationary model or vector-like fermions in a gauged $U(1)$ extension of the SM mentioned above. Then the scalar froze out at some temperature $T_{f}$ and eventually it dominates the energy budget of the Universe at temperature $T_{\rm dom}$. If the freeze out happens while the scalar particles were still relativistic we obtain the maximal $\phi$ abundance Bezrukov:2009th ; Kolb:1990vq ,
@@
The effective degrees of freedom related to entropy is denoted by $g_{*S}$ which we take to be equal to $g_{*}=106.75$ at high temperatures for simplicity in our analysis. After freeze-out, $\phi$ eventually becomes non-relativistic after which its energy density red-shifts as matter and dominates over the radiation density starting from temperature Giudice:1999fb ; Kolb:1990vq ,
@@
The $\phi$-dominated phase starts at temperature $T_{\rm dom}$ and ends at some temperature decay temperature $T_{\phi}$; we always assume $T_{\phi}<T_{\rm dom}$ such that we get a period of time in early Universe of scalar era. The dynamics of $\phi$ and its decay products into RHNs and radiation can be obtained by solving the relevant Boltzmann equations which we will describe in Sec. III.2.
We start with Eq. (20) and (21), the partial decay widths of $\phi$ to lightest RHNs and radiation and consider only the two lightest RHNs $N_{1},N_{2}$. The branching ratios are given by (22). As long as mass of the scalar $M_{\phi}>2M_{1,2}$, the scalar can decay to $N_{1}$ and $N_{2}$.
@@
The RHNs $N_{1},N_{2}$ further decay to lepton-Higgs pairs with the thermally averaged decay widths,
$\Gamma_{N_{i}}=\Gamma_{N_{i}}^{rf}\,\frac{K_{1}(z)}{K_{2}(z)}\sim\mathcal{H}(M_{i})\,K\,\frac{K_{1}(z)}{K_{2}(z)},$ (26)
where $\mathcal{H}(M_{i})=\sqrt{\frac{8\pi^{3}g_{\ast}}{90}}\frac{M_{i}^{2}}{M_{\text{pl}}}$ is the Hubble parameter with $M_{\text{pl}}=1.22\times 10^{19}$ GeV. $K_{1}(z)$,$K_{2}(z)$ are the modified bessel functions of first and second kind and $z=M_{1}/T$.
@@
The decay temperatures of the scalar or the RHNs are calculated assuming that the energy density of scalar/RHN is converted instantaneously into radiation Hahn-Woernle:2008tsk ,
$T_{\phi}=\left(\frac{90}{8\pi^{3}g_{\ast}(T_{\phi})}\right)^{\frac{1}{4}}\sqrt{\Gamma_{\phi}M_{\text{pl}}}$ (27)
$T_{N_{i}}=\left(\frac{90}{8\pi^{3}g_{\ast}(T_{N_{i}})}\right)^{\frac{1}{4}}\sqrt{\Gamma_{N_{i}}^{rf}M_{\text{pl}}}=M_{i}\sqrt{K}$ (28)
where $\Gamma_{\phi}\sim\Gamma_{\phi\rightarrow N_{1}N_{1}}+\Gamma_{\phi\rightarrow N_{2}N_{2}}+\Gamma_{\phi\rightarrow R}$ is the total decay width of $\phi$ and $g_{\ast}(T)$ is the relativistic degree of freedom at the corresponding temperature $T$ which is taken to be $\sim 106.75$. Using Eq. (20) (21) in Eq. (27) and considering Eq. (25), we can write the condition $T_{\phi}<T_{\rm dom}$ in terms of couplings as,
$y_{N_{i}}^{2}+y_{R}^{2}\lesssim\frac{M_{\phi}}{M_{\rm pl}}$ (29)
where $i=1,2$ and we have considered $M_{1},M_{2}\ll M_{\phi}$. We can expect that the latest radiation dominated era starts at either $T_{\phi}$ or $T_{N_{1}}$ depending on which one is lower. Notice that reheating temperature $T_{\text{RH}}$ after inflation now need not be larger than $M_{1}$ as in the thermal case. Rather $T_{\rm RH}$ only needs to be roughly larger than $T_{\rm dom}$ now, since RHNs are produced dominantly at $T_{\phi}<T_{\rm dom}$ from $\phi$ decay, regardless of their masses.
@@
In Table 1, we summarize all the necessary conversion formulas for the parameters used in our analysis in terms of the couplings and masses of the scalar and RHNs.
@@
RHN decay width in the rest frame $\Gamma_{N_{i}}^{rf}$ $\frac{|\lambda|^{2}M_{i}}{8\pi}$
Scalar decay width to RHNs $\Gamma_{\phi\rightarrow N_{i}N_{i}}$ $\frac{y_{N_{i}}^{2}M_{\phi}}{16\pi}\sqrt{1-\frac{4M_{i}^{2}}{M_{\phi}^{2}}}$
Scalar decay width to radiation $\Gamma_{\phi\rightarrow R}$ $\frac{y_{R}^{2}\,M_{\phi}}{8\pi}$
Total decay width of scalar $\Gamma_{\phi}$ $\Gamma_{\phi\rightarrow N_{1}N_{1}}+\Gamma_{\phi\rightarrow N_{2}N_{2}}+\Gamma_{\phi\rightarrow R}$
Scalar decay temperature $T_{\phi}$ $\left(\frac{90}{8\pi^{3}g_{\ast}(T_{\phi})}\right)^{\frac{1}{4}}\sqrt{\Gamma_{\phi}M_{\text{pl}}}$
$\sim 0.05\sqrt{y_{N_{1}}^{2}+y_{R}^{2}}\left(\frac{100}{g_{*}(T_{\phi})}\right)^{\frac{1}{4}}\sqrt{M_{\phi}M_{\rm pl}}$
Washout parameter $K$ $\frac{\tilde{m}_{1}}{m_{*}}$=$\frac{\tilde{m}_{1}}{1.1\times 10^{-3}\,\text{eV}}=\frac{|\lambda|^{2}_{11}v^{2}/M_{1}}{1.1\times 10^{-3}\,\text{eV}}$
@@
Table 1: Definition of the relevant parameters used for our analysis in terms of couplings and masses of the scalar and RHNs. Here $v=174$ GeV is the electrweak VEV, $\lambda$ is the Yukawa coupling of RHNs with active neutrinos via SM Higgs. Here we also considered resonant scenario, i.e. $y_{N_{1}}\sim y_{N_{2}}\ll y_{N_{3}}$.

### v2

@@ II.2 CP Asymmetry
The CP asymmetry $\epsilon_{1}$ is determined by the interference between the tree level and 1-loop diagrams  Flanz:1994yx ; Covi:1996wh and can be calculated from $lH\rightarrow lH$ scattering processes Buchmuller:1997yu . The relevant Feynmann diagrams giving tree level, vertex, and self-energy contributions are shown in Fig. 1.
Figure 1: Feynmann diagrams of $N\rightarrow lH$ decay process at the (a) tree-level and (b)-(c) 1-loop level. (b) represents the vertex contribution and (c) represents the self-energy contribution.
For RHN mass hierarchy $M_{1}<M_{2}<M_{3}$, the CP parameter is obtained as,
$|\epsilon_{1}|=\sum_{i\neq 1}\frac{3}{16\pi}\frac{M_{1}}{M_{i}}\frac{\text{Im}\left[\left(\lambda\lambda^{\dagger}\right)_{1i}^{2}\right]}{|\lambda\lambda^{\dagger}|_{11}}=\frac{3}{16\pi}\frac{M_{1}}{v^{2}}\frac{\sum_{i}m_{i}^{2}\text{Im}\left[R_{1i}^{2}\right]}{\sum_{j}m_{j}|R_{1j}|^{2}}<\epsilon_{1}^{\rm max}$ (18)
where the maximal asymmetry is given by Buchmuller:2003gz ; Hahn-Woernle:2008tsk ; Buchmuller:2004tu ; Davidson:2002qv ; Hambye:2003rt ,
@@
Note that for small values of $\tilde{m}_{1}$ we require small values of $|R_{1i}|^{2}$ as seen from Eq. (12). However, that does not necessarily make $|\epsilon_{1}|$ tiny as it depends on the ratio of off-diagonal elements of $R$ as seen in Eq. (18). It can be verified that it is possible to achieve maximal CP asymmetry even if $\tilde{m}_{1}\rightarrow 0$, implying $K\rightarrow 0$ (see Appendix A). For our analysis, we will consider the range $K\in[10^{-20},10^{4}]$ while keeping $K_{2},K_{3}>1$.
@@ II.4 Resonant Leptogenesis
Thermal leptogenesis operates at high temperatures, which in turn pushes the inflationary reheating temperature $T_{\rm RH}$ to a high scale, as shown in Eq. (20) and (21). Such high scales involving RHN masses $M_{i}\gtrsim 10^{9}$ GeV are difficult to test in the lab. Also, a high neutrino mass conflicts with the Electroweak naturalness condition which requires RHN masses $M_{i}\lesssim 10^{7}$ GeV Vissani:1997ys ; Clarke:2015gwa ; Datta:2022tab . A lower reheating temperature from inflation $T_{\rm RH}$ is motivated in different inflationary scenarios like Drees:2021wgd ; Ghoshal:2022zwu . Even in some SUGRA models Kawasaki:2004qu , reheating temperature is found to be as low as $10^{6}$ GeV which contradicts a high $T_{\text{RH}}$, favoring low-scale-leptogenesis. Motivated by these reasons, and for lab testability, the scale of leptogenesis can be brought down by several orders in the so-called resonant regime which focuses on nearly degenerate heavy neutrinos leading to pronounced CP violation in their decays. It can be shown that if the mass difference between the two lightest right-handed neutrinos is of the order of the decay widths of them, i.e. $M_{2}-M_{1}\sim\Gamma_{N_{2}}^{rf}$ the CP parameter $\epsilon_{1}$ is enhanced via the self-energy contribution Pilaftsis:1997dr ; Hambye:2003rt in the loop level diagrams shown in Fig 1.
@@
where, $\epsilon_{1}^{\rm max}$ is the maximal asymmetry in non-resonant case given by Eq. (19). It was shown that within the resonant regime, the lowest RHN mass $M_{1}$ can be as low as 1 TeV Pilaftsis:1997jf . In Pilaftsis:2005rv this scale is shown to reach upto the electroweak symmetry breaking scale $\sim 100$ GeV, below which sphaleron processes are suppressed and the lepton asymmetry can not be effectively converted to baryon asymmetry. From here onwards we will consider resonant leptogenesis with $M_{1}\gtrsim 100$ GeV for our analysis. Note that in some alternate leptogenesis scenarios, e.g. via neutrino oscillations, BAU can be generated for even for RHN mass values $\sim 100$ MeV Klaric:2020phc .
III BLB-L Symmetry Breaking and Non-thermal Leptogenesis
In this section, we construct a global $U(1)_{B-L}$ neutrino mass model inspired by Fong:2023egk , that includes all the essential components for non-thermal leptogenesis. In non-thermal leptogenesis, RHNs are produced from inflaton decay Asaka:1999yd ; Giudice:2003jh ; Hahn-Woernle:2008tsk ; Zhang:2023oyo or from a generic scalar decay Datta:2022tab ; Nemevsek:2023yjl . Such a scenario can lead to an excess initial abundance of RHNs compared to their thermal abundance. As a result, the final efficiency factor, $\kappa_{f}$ can be enhanced by a factor 100 which relaxes the Davidson-Ibarra bound on RHN mass given in Eq. 20. In this paper, the model extends the SM to include right handed neutrinos $N_{i}$ where $i\in\{1,2,3\}$, and an additional scalar boson denoted as $\sigma$, with BLB-L charge of +2+2. This complex scalar is responsible for generating the heavy Majorana masses of RHNs after spontaneously breaking the $U(1)_{B-L}$ symmetry. The Lagrangian is as follow,
@@ III BLB-L Symmetry Breaking and Non-thermal Leptogenesis
$V(H,\sigma)$ $=\lambda_{H}(|H|^{2}-v^{2})^{2}+\lambda_{\sigma}(|\sigma|^{2}-v_{B-L}^{2})^{2}+\lambda_{\sigma H}|\sigma|^{2}|H|^{2}\,$, (23)
where $L$ is the leptonic doublet, $H$ denotes the Higgs doublet, and $\tilde{H}=i\sigma_{2}H^{*}$ as in Eq. (2). The symbols $\lambda$, $y_{N}$, $\lambda_{H}$, $\lambda_{\sigma H}$, and $\lambda_{\sigma}$ encompass the dimensionless coupling constants in the theory. The Yukawa coupling matrices $\lambda$ and $y_{N}=\text{diag}[\![y_{N_{i}}]\!]$ parameterize the strengths of the SM Higgs and the scalar couplings to the RHNs. Here $v_{B-L}$ is the BLB-L symmetry-breaking scale. These parameters, as we will see, play key roles in determining the GW spectral shapes which we explore in details. Around the BL{B-L} breaking vacuum, $\sigma$ can be parameterized as,
$\sigma$ $=$ $(v_{B-L}+\rho)e^{i\phi/v_{B-L}}\,$, (24)
where $\phi$ is the Nambu-Goldstone (NG) boson called majoron. After BLB-L symmetry is broken, the $N_{i}$ acquires the Majorana mass $M_{i}=y_{N_{i}}v_{B-L}$ where the extra phase is rotated away by redefining all fermions as $f\rightarrow f^{\prime}=fe^{iq_{f}^{B-L}\phi/2v_{B-L}}$, where $q_{f}^{B-L}$ is the BLB-L charge of the fermion $f$. The radial component $\rho$ acquires a mass $M_{\rho}=\sqrt{8\lambda_{\sigma}}v_{B-L}$ in the vanishing portal coupling limit, $\lambda_{\sigma H}=0$. The majoron $\phi$ couples to the RHNs through its kinetic term which gives,
$\mathcal{L}\supset-\frac{\partial_{\mu}\phi}{2v_{B-L}}J^{\mu}_{B-L}=\frac{\phi}{2v_{B-L}}\partial_{\mu}J^{\mu}_{B-L},$ (25)
where the BLB-L current is $J^{\mu}_{B-L}=\sum_{f}q^{B-L}_{f}\overline{f}\gamma^{\mu}f$. The second part of the equation is obtained by carrying out the integration by parts in the action and ignoring the surface term. This equation means that the majoron couples to BLB-L violating currents, which in the RHN sector is the Majorana mass term of the RHNs. Hence, we obtain the majoron-RHN coupling,
$\mathcal{L}\supset\frac{i\phi}{v_{B-L}}M_{i}\overline{N^{c}_{i}}N_{i}.$ (26)
At this point, the majoron is massless. To give it mass, we consider a asymptotically free hidden gauge sector with confinement scale $\Lambda<v_{B-L}$, and make it anomalous under the $U(1)_{B-L}$, by adding fermions which are vector-like under the SM but chiral under the $U(1)_{B-L}$ and charged under the hidden group. This group becomes non-perturbative at the scale $\Lambda$, and the majoron couples to the new gauge bosons as follows,
$\mathcal{L}\supset\frac{\phi}{v_{B-L}}\frac{g_{X}^{2}n}{16\pi^{2}}G\tilde{G}$, (27)
where $g_{X}$ is the new gauge coupling, $n$ is the anomaly coefficient and $\tilde{G}$ is the dual of the field strength tensor $G$ of the new gauge group. This term breaks the shift symmetry of $\phi$ and generates a periodic potential of $\phi$ via instantons,
$V(\phi)=\Lambda^{4}\left(1-\cos\frac{\phi}{v_{B-L}}\right),$ (28)
The mass of the majoron can be found by,
$M_{\phi}^{2}=\frac{\partial^{2}V}{\partial\phi^{2}}\Bigg|_{\phi=0}=\frac{\Lambda^{4}}{v_{B-L}^{2}}.$ (29)
We will assume that $M_{\rho}\sim v_{B-L}\gg T_{\rm RH}$ such that effects of $\rho$ can be ignored. We will also assume the majoron mass to always exceed twice the lightest RHN mass, $M_{\phi}>2M_{1}$. Moreover we consider a resonant leptogenesis scenario with $M_{1}\sim M_{2}\ll M_{3}$. The resonance effect makes sure that enough asymmetry is generated for all $M_{1}$ values starting from 100 GeV, to explain the BAU. For simplicity we take $M_{3}>M_{\phi}$. The decay width of $\phi$ to RH neutrinos is,
$\Gamma_{\phi\to N_{i}N_{i}}$ $=$ $\frac{M_{\phi}M_{i}^{2}}{16\pi v_{B-L}^{2}}\sqrt{1-\frac{4M_{i}^{2}}{M_{\phi}^{2}}}\,$, (30)
where $i=1,2$ represents the 2 lightest RH neutrinos. The majoron decay to the hidden sector through the coupling in Eq. (27) is loop suppressed or forbidden if the hidden sector particles are heavier than $\phi$.
@@
We consider the parameter space of the majoron where it dominates the energy budget of the Universe before decaying, leading to a phase of matter-dominated pre-BBN era. The RHNs produced from majoron decay further dominate the energy density. The decay of majoron and RHNs inject considerable amounts of entropy into the Universe. This dampens the primordial gravitational waves of inflatinary origin that re-enter the horizon during the early matter domination epochs. We show such signature of low-scale leptogenesis in primordial gravitational waves can be detected in various GW experiments such as BBO, DECIGO, $\mu$-ARES, ET, LISA, CE, etc.
@@ III.1 Set Up
After inflationary reheating ends at temperature $T_{\rm RH}$, the majorons are produced either a) thermally via suppressed scatterings, e.g. $lH\rightarrow\phi N$ or b) non-thermally from inflaton and $\rho$-decay or through misalignment (coherrent oscillations) starting around $T_{\rm osc}\sim\sqrt{M_{\phi}M_{\rm Pl}}$. We make the coupling $y_{N}$ small such that $\phi$ is long-lived, hence no production through $N$ inverse decay is possible. The misalignment mechanism gives an energy density of the majoron,
$\rho_{\phi}(T_{\rm osc})\sim\frac{1}{2}M_{\phi}^{2}\phi_{\rm init}^{2}$ (31)
where $\phi_{\rm init}$ is the initial value of $\phi$ at $T_{\rm osc}$. We consider that $\phi$ had an abundance equal to the thermal abundance at temperature $T\sim M_{\phi}$ after the end of inflation, i.e. $\rho_{\phi}(M_{\phi})=M_{\phi}n_{\phi}^{eq.}$. We can adjust $\phi_{\rm init}$ to achieve this result. In thermal case, if scatterings freeze-out while the majoron particles were still relativistic we obtain the maximal $\phi$ abundance Bezrukov:2009th ; Kolb:1990vq ,
@@
The effective number of degrees of freedom related to entropy is denoted by $g_{*S}$ which we take to be equal to $g_{*}=106.75$ at high temperatures for simplicity in our analysis. In both the cases: misalignment or thermal, $\phi$ is non-relativistic and frozen-out at $T\sim M_{\phi}$, after which its energy density red-shifts as matter and dominates over the radiation density starting from temperature Giudice:1999fb ; Kolb:1990vq ,
@@
The $\phi$-dominated phase starts at temperature $T_{\rm dom}$ and ends at decay temperature $T_{\phi}$, calculated by considering the sudden decay assumption,
$T_{\phi}=\left(\frac{90}{8\pi^{3}g_{\ast}(T_{\phi})}\right)^{\frac{1}{4}}\sqrt{\Gamma_{\phi}M_{\text{pl}}}.$ (35)
@@
We always assume $T_{\phi}<T_{\rm dom}$ such that we get a period of majoron domination in the early Universe. The dynamics of $\phi$, RHNs and radiation can be obtained by solving the relevant Boltzmann equations which we will describe in Sec. III.2.
Now let us look at the aftermath of $\phi$ decay. We start with Eq. (30), the decay widths of $\phi$ to lightest RHNs. As long as the mass of the majoron $M_{\phi}>2M_{1,2}$, it can decay to $N_{1,2}$. The majorons we consider are essentially heavy ($10^{5}-10^{15}$ GeV) and non-relativistic in post-inflationary scenario and their decay is in their rest frame. Hence, RHNs resulting from this can reasonably be monochromatic, where each RHN will approximately have energy $E_{N}\sim M_{\phi}/2$ at production. We are concerned only about the decay dynamics of $N_{1}$ Engelhard:2007kf . If $M_{\phi}\gg M_{1}$, the RHNs are relativistic when they are produced at temperatures around $T_{\phi}$. The energy of these relativistic RHNs, $E^{\rm rel}_{N}(T)$ redshifts with the expansion of the Universe such that $E^{\rm rel}_{N}(T)/T=E_{N}/T_{\phi}$. This gives,
$E^{\rm rel}_{N}(z)\sim E_{N}\frac{T}{T_{\phi}}\sim\frac{M_{\phi}}{2\,T_{\phi}}\frac{M_{1}}{z},$ (36)
where $z=M_{1}/T$. The RHNs then become non-relativistic when their energy becomes equivalent to their mass, i.e. at temperature Zhang:2023oyo ,
$T_{\rm NR}\sim T_{\phi}\frac{2M_{1}}{M_{\phi}}.$ (37)
Hence, the energy of each RHN after its production can be written as,
$E_{N}(z)\sim\begin{cases}\frac{M_{\phi}}{2}&\text{for $T>T_{\phi}$}\\ \frac{M_{\phi}}{2\,T_{\phi}}\frac{M_{1}}{z}&\text{for $T_{\phi}\geq T>T_{\rm NR}$}\\ M_{1}&\text{for $T\leq T_{\rm NR}$}\end{cases}$ (38)
@@
The RHNs further decay to lepton-Higgs pairs with decay widths Cataldi:2024pgt ,
$\Gamma_{N_{1}}=\Gamma_{N_{1}}^{rf}\,\frac{M_{1}}{E_{N}(z)}\sim\mathcal{H}(M_{1})\,K\,\frac{M_{1}}{E_{N}(z)},$ (39)
where $\mathcal{H}(M_{1})=\sqrt{\frac{8\pi^{3}g_{\ast}}{90}}\frac{M_{1}^{2}}{M_{\text{pl}}}$ is the Hubble parameter with $M_{\text{pl}}=1.22\times 10^{19}$ GeV. Note that the decay rates get suppressed by the Lorentz boost or time dilation factor $\gamma_{N}=E_{N}(z)/M_{1}$.
@@
The decay temperatures of the RHNs is calculated assuming that the energy density of RHN is converted instantaneously into radiation Hahn-Woernle:2008tsk ,
$T_{N_{1}}=\left(\frac{90}{8\pi^{3}g_{\ast}(T_{N_{1}})}\right)^{\frac{1}{4}}\sqrt{\Gamma_{N_{1}}^{rf}M_{\text{pl}}}=M_{1}\sqrt{K}$ (40)
where $g_{\ast}(T)$ is the relativistic degrees of freedom at the corresponding temperature $T$ which is taken to be $\sim 106.75$. The final baryon asymmetry from RHN decay can be written as,
$Y_{B}=\frac{n_{B}-n_{\overline{B}}}{s}=\frac{a_{\rm sph}\epsilon}{s}n_{N}$ (41)
where $s=g_{*}(2\pi^{2}/45)T^{3}$ is the entropy density and $n_{B},n_{\overline{B}},n_{N}$ are the number densities of baryons, anti-baryons and RHNs respectively. $Y_{B}$ is related to $\eta_{B}$ given in Eq. (6) as, $Y_{B}=(n_{\gamma 0}/s_{0})\eta_{B}\simeq\eta_{B}/7.04$ Davidson:2008bu , with $s_{0}$, $n_{\gamma 0}$ being the current entropy and photon number densities respectively. The factor $a_{\rm sph}=28/79$ is the efficiency of the sphaleron processes.
Using Eq. (30) in Eq. (35) and considering Eq. (34), we can write the condition $T_{\phi}<T_{\rm dom}$ in terms of couplings as,
$y_{N_{1}}\lesssim 0.2\sqrt{\frac{M_{\phi}}{M_{\rm pl}}}.$ (42)
@@
In Table 1, we summarize all the necessary conversion formulas for the parameters used in our analysis in terms of the couplings and masses of $\phi$ and RHNs. We consider $M_{\phi},T_{\phi},M_{1},K$ and $T_{\rm RH}$ as input parameters for our analysis and $v_{B-L},y_{N_{1}},|\lambda|^{2}_{11},T_{N_{1}},\Gamma_{\phi},\Gamma_{N_{1}}^{rf}$ are derived from them.
@@
RHN decay width in the rest frame $\Gamma_{N_{1}}^{rf}$ $\frac{|\lambda|^{2}_{11}M_{1}}{8\pi}$
Majoron decay width to RHNs $\Gamma_{\phi\rightarrow N_{i}N_{i}}$ $\frac{y_{N_{i}}^{2}M_{\phi}}{16\pi}\sqrt{1-\frac{4M_{i}^{2}}{M_{\phi}^{2}}}$
Total decay width of majoron $\Gamma_{\phi}$ $\Gamma_{\phi\rightarrow N_{1}N_{1}}+\Gamma_{\phi\rightarrow N_{2}N_{2}}$
Majoron decay temperature $T_{\phi}$ $\left(\frac{90}{8\pi^{3}g_{\ast}(T_{\phi})}\right)^{\frac{1}{4}}\sqrt{\Gamma_{\phi}M_{\text{pl}}}$
$\sim 0.05\,y_{N_{1}}\left(\frac{100}{g_{*}(T_{\phi})}\right)^{\frac{1}{4}}\sqrt{M_{\phi}M_{\rm pl}}$
Washout parameter $K$ $\frac{\tilde{m}_{1}}{m_{*}}$$\frac{\tilde{m}_{1}}{1.1\times 10^{-3}\,\text{eV}}=\frac{|\lambda|^{2}_{11}v^{2}/M_{1}}{1.1\times 10^{-3}\,\text{eV}}$
@@
Table 1: Definition of the key parameters used for our analysis in terms of couplings and masses of the majoron and RHNs. Here $v=174$ GeV is the electroweak VEV, and $\lambda$ is the Yukawa coupling of RHNs with active neutrinos via SM Higgs. Here we also considered $y_{N_{2}}\sim y_{N_{1}}$.